OpenStudy (anonymous):

. A basket contains 11 pieces of fruit: 4 apples, 5 oranges, and 2 bananas. Jonas takes a piece of fruit at random from the basket, and then Beth takes a piece at random. What is the probability that Jonas will get an orange and Beth will get an apple? A. mc007-1.jpg B. mc007-2.jpg C. mc007-3.jpg D. mc007-4.jpg

OpenStudy (anonymous):

someone please help this is timed

OpenStudy (nurali):

Total outcomes = 11 favorable outcomes = 5 Probability = Number of favorable outcomes / Total number of outcomes Probability that Jonas will get an orange = 5/11 Total outcomes = 10 (because 1 fruit is already taken by Jonas) favorable outcomes = 4 Probability Beth will get an apple = 4/10 = 2/5 multiply both probability that Jonas will get an orange and Beth will get an apple = 5/11 * 2/5 = 2/11

OpenStudy (kendricklamar2014):

Welcome to Open Study, this is History and you'r question should be in the Mathematics subject. The Jpg's you sent us, we cant see the. So we can see them, put them on a reply, and click the "Attach File" button. :)